1) Solve the following:

Q                                       Total Utility

0                                              0         

1    60

2                                              130

3                                              190

4                                              230

5                                              250

Calculate her Marginal Utility at Q 2?

2) Marginal Utility at Q 4?

3) Average Utility at Q 3?

4) Average Utility at Q 5?
Answer #1

Answer : Table showing information :

Quantity Total utility Marginal utility Average utility = Utility / Quantity
0 0 - -
1 60 60 60
2 130 70 65
3 190 60 63.3
4 230 40 57.5
5 250 20 50

1. Marginal utility at Q 2= 130-60= 70utilies

2.Marginal utility at Q4 = 230-190=40utilies

3.Average utility at Q3 = 190/3 = 63.3utilies

4.Average utility at Q5 = 250/5 =50utilies